# A Bounding Box is Worth One Token: Interleaving Layout and Text in a Large Language Model for Document Understanding

### v1

@@ Abstract
Recently, many studies have demonstrated that exclusively incorporating OCR-derived text and spatial layouts with large language models (LLMs) can be highly effective for document understanding tasks. However, existing methods that integrate spatial layouts with text have limitations, such as producing overly long text sequences or failing to fully leverage the autoregressive traits of LLMs. In this work, we introduce Interleaving Layout and Text in a Large Language Model (LayTextLLM) for document understanding. In particular, LayTextLLM projects each bounding box to a single embedding and interleaves it with text, efficiently avoiding long sequence issues while leveraging autoregressive traits of LLMs. LayTextLLM not only streamlines the interaction of layout and textual data but also shows enhanced performance in Key Information Extraction (KIE) and Visual Question Answering (VQA). Comprehensive benchmark evaluations reveal significant improvements, with a 27.0% increase on KIE tasks and 24.1% on VQA tasks compared to previous state-of-the-art document understanding MLLMs, as well as a 15.5% improvement over other SOTA OCR-based LLMs on KIE tasks.
@@ 4.2 Implementation Detail
We have configured the model with three versions of LayTextLLM for a side-by-side comparison under different settings. Aligned with Luo et al. [27], the first version, LayTextLLMzero, is trained exclusively with DDD and Layout-aware SFT data. Building upon this, and in alignment with the setting of Liu et al. [12], we introduce the DocVQA, InfoVQA, and ChartQA training sets to the dataset pool for our second version, termed LayTextLLMvqa. Finally, we incorporate a comprehensive suite of KIE datasets—FUNSD, CORD, POIE, SROIE, and VisualMRC—as described by Wang et al. [25], creating our most extensive version, LayTextLLMall. Note that all versions are based on the same pre-trained LayTextLLM weight.
@@ 4.4 Evaluation Metrics
LayTextLLMzero (Ours) 71.8 33.8 52.8 49.4 86.7 66.1 67.4
LayTextLLMvqa (Ours) 77.4 66.1 71.8 48.9 74.6 70.6 64.7
@@
DocLLM [25] 69.5 264.1 166.8 51.8 67.4 91.9 70.3
LayTextLLMzero (Ours) 65.4 260.7 163.0 48.6 74.5 86.4 69.8
LayTextLLMvqa (Ours) 75.7 260.2 168.0 52.7 70.9 78.6 67.4
LayTextLLMall (Ours) 77.3 295.9 186.6 64.2 96.5 95.8 85.8
@@
LayTextLLMzero (Ours) 65.4 36.2 50.8 71.0 66.9 89.2 75.7
LayTextLLMall (Ours) 77.3 41.7 59.5 81.3 82.6 96.2 86.7
@@ 4.5 Quantitative Results
The experimental results shown in Tab. 1 demonstrate the outstanding performance of the LayTextLLM series across various tasks. Note that the results for ChartQA are reported in Appendix E due to concerns about fairness in comparison, as the dataset does not include OCR-derived results and we used in-house OCR tools instead. Firstly, LayTextLLMzero significantly outperforms previous SOTA OCR-free methods, such as TextMonkey [12], in zero-shot capabilities, even when these methods use the training set of the dataset. For example, in the DocVQA and InfoVQA datasets, LayTextLLMzero achieves accuracies of 71.8% and 33.8%, respectively, which are markedly higher than existing OCR-free methods such as TextMonkey and InternLM-XComposer2. When fine-tuned with corresponding datasets, LayTextLLM shows even greater performance improvements, particularly in document-oriented VQA datasets. Specifically, its accuracies on DocVQA and InfoVQA increase to 77.4% and 66.1%, respectively, demonstrating the model’s strong ability to leverage task-specific data. Additionally, LayTextLLMzero excels in KIE datasets, particularly on the SROIE and POIE datasets, achieving accuracies of 86.7% and 66.1%, respectively. These results significantly surpass those of previous SOTA OCR-free model (i.e., TextMonkey+) by margins of 40.5% and 34.1%, respectively. This significant performance gain is likely due to these datasets containing low-resolution images that are too blurred for current MLLMs to extract visual features, whereas LayTextLLM shows robustness in such challenging scenarios.
@@ 4.6 Analysis
Ablations To better assess the utility of each component in LayTextLLM, an ablation study was conducted, the results of which are presented in Tab. 4. Detailed information on the training setup for all variants is provided in Appendix B. The results clearly show that incorporating interleaved spatial layouts and texts significantly enhances the performance, evidenced by a 2.8% improvement in VQA and a 6.6% increase in KIE compared to the plain version, indicating that SLP is a critical component. Furthermore, enabling P-LoRA results in a modest performance increase of 0.8% in both VQA and KIE tasks. Finally, the enabling of LNTP+SSFT leads to a significant improvement in VQA, with an increase of 3.8% though with a slight decrease with KIE tasks.
Sequence Length Tab. 5 presents statistics on the average input sequence length across different datasets. Intriguingly, despite interleaving bounding box tokens, LayTextLLM consistently exhibits the shortest sequence length in three out of four datasets, even surpassing DocLLM, which is counterintuitive. We attribute this to the tokenizer mechanism. For example, using tokenizer.encode(), a single word from the OCR engine, like “International” is encoded into a single ID $[4623]$. Conversely, when the entire OCR output is processed as one sequence, such as “… CPC,International,Inc…”, the word “International” is split into two IDs $[17579,1288]$, corresponding to “Intern” and “ational” respectively. This type of case occurs frequently, more discussion in Appendix C.
@@ Appendix E Results of ChartQA
As shown in Fig. 7, the question-answer pairs in ChartQA [49] tend to involve the visual cues for reasoning. However, with only text and layout information as input, the proposed LayTextLLM inevitably have difficulties in reasoning visual-related information. Thus, on the ChartQA dataset, LayTextLLM can hardly achieve better performance than previous methods that include visual inputs. Although the visual information is not used in LayTextLLM, it can still exhibit better zero-shot ability than UniDoc [10]. After incorporating the training set of ChartQA, the performance of LayTextLLM can be boosted to 35.7%. Considering the importance of visual cues in ChartQA-like tasks, we will try to involve the visual information into LayTextLLM in future work.
@@
LayTextLLMzero (Ours) 21.4
LayTextLLMvqa (Ours) 29.8
LayTextLLMall (Ours) 35.7

### v2

# LayTextLLM: A Bounding Box is Worth One Token - Interleaving Layout and Text in a Large Language Model for Document Understanding

@@ Abstract
Recently, many studies have demonstrated that exclusively incorporating OCR-derived text and spatial layouts with large language models (LLMs) can be highly effective for document understanding tasks. However, existing methods that integrate spatial layouts with text have limitations, such as producing overly long text sequences or failing to fully leverage the autoregressive traits of LLMs. In this work, we introduce Interleaving Layout and Text in a Large Language Model (LayTextLLM) for document understanding. In particular, LayTextLLM projects each bounding box to a single embedding and interleaves it with text, efficiently avoiding long sequence issues while leveraging autoregressive traits of LLMs. LayTextLLM not only streamlines the interaction of layout and textual data but also shows enhanced performance in Key Information Extraction (KIE) and Visual Question Answering (VQA). Comprehensive benchmark evaluations reveal significant improvements, with a 27.2% increase on KIE tasks and 12.0% on VQA tasks compared to previous state-of-the-art document understanding MLLMs, as well as a 15.1% improvement over other SOTA OCR-based LLMs on KIE tasks. Furthermore, we found that the spatial layout can be decoded back into coordinates, and inference requiring output bounding box coordinates can further alleviate the hallucination problem. All resources are available at https://github.com/LayTextLLM/LayTextLLM.
@@ 4.2 Implementation Detail
We have configured the model with three versions of LayTextLLM for a side-by-side comparison under different settings. Aligned with Luo et al. [27], the first version, LayTextLLMzero, is trained exclusively with DDD and Layout-aware SFT data. Building upon this, and in alignment with the setting of Liu et al. [12], we introduce the DocVQA and InfoVQA training sets to the dataset pool for our second version, termed LayTextLLMvqa. Finally, we incorporate a comprehensive suite of KIE datasets—FUNSD, CORD, POIE, SROIE, ChartQA, and VisualMRC—as described by Wang et al. [25], creating our most extensive version, LayTextLLMall. Note that all versions are based on the same pre-trained LayTextLLM weight.
@@ 4.4 Evaluation Metrics
LayTextLLMzero (Ours) 72.1 35.7 53.9 47.5 86.4 68.9 67.6
LayTextLLMvqa (Ours) 77.2 42.1 59.7 48.8 75.7 70.6 65.0
@@
DocLLM [25] 69.5 264.1 166.8 51.8 67.6 91.9 70.3
LayTextLLMzero (Ours) 65.5 200.2 132.9 47.2 77.2 83.7 69.4
LayTextLLMvqa (Ours) 75.6 179.5 127.6 52.6 70.7 79.3 67.5
LayTextLLMall (Ours) 77.2 277.8 177.6 64.0 96.5 95.8 85.4
@@
LayTextLLMzero (Ours) 65.5 37.4 51.5 72.0 45.5 82.0 66.5
LayTextLLMall (Ours) 77.2 41.7 59.5 81.0 82.5 96.1 86.5
@@ 4.5 Quantitative Results
The experimental results shown in Tab. 1 demonstrate the outstanding performance of the LayTextLLM series across various tasks. Note that the results for ChartQA are reported in Appendix E due to concerns about fairness in comparison, as the dataset does not include OCR-derived results and we used in-house OCR tools instead. Firstly, LayTextLLMzero significantly outperforms previous SOTA OCR-free methods, such as TextMonkey [12], in zero-shot capabilities, even when these methods use the training set of the dataset. For example, in the DocVQA and InfoVQA datasets, LayTextLLMzero achieves accuracies of 72.1% and 35.7%, respectively, which are markedly higher than existing OCR-free methods such as TextMonkey and InternLM-XComposer2. When fine-tuned with corresponding datasets, LayTextLLM shows even greater performance improvements, particularly in document-oriented VQA datasets. Specifically, its accuracies on DocVQA and InfoVQA increase to 77.2% and 42.1%, respectively, demonstrating the model’s strong ability to leverage task-specific data. Additionally, LayTextLLMzero excels in KIE datasets, particularly on the SROIE and POIE datasets, achieving accuracies of 86.4% and 68.9%, respectively. These results significantly surpass those of previous SOTA OCR-free model (i.e., TextMonkey+) by margins of 40.5% and 34.1%, respectively. This significant performance gain is likely due to these datasets containing low-resolution images that are too blurred for current MLLMs to extract visual features, whereas LayTextLLM shows robustness in such challenging scenarios.
@@ 4.6 Analysis
Ablations To better assess the utility of each component in LayTextLLM, an ablation study was conducted, the results of which are presented in Tab. 4. Detailed information on the training setup for all variants is provided in Appendix B. The results clearly show that incorporating interleaved spatial layouts and texts significantly enhances the performance, evidenced by a 2.8% improvement in VQA and a 6.6% increase in KIE compared to the plain version, indicating that SLP is a critical component. Furthermore, enabling P-LoRA results in a modest performance increase of 0.8% in both VQA and KIE tasks. Finally, the enabling of LNTP+SSFT leads to a significant improvement in VQA and KIE tasks, with an increase of 3.6% and 0.1%.
Sequence Length Tab. 5 presents statistics on the average input sequence length across different datasets. Intriguingly, despite interleaving bounding box tokens, LayTextLLM consistently exhibits the shortest sequence length in three out of four datasets, even surpassing DocLLM, which is counterintuitive. We attribute this to the tokenizer mechanism. For example, using tokenizer.encode(), a single word from the OCR engine, like “International” is encoded into a single ID $[4623]$. Conversely, when the entire OCR output is processed as one sequence, such as “… CPC,International,Inc…”, the word “International” is split into two IDs $[17579,1288]$, corresponding to “Intern” and “ational” respectively. This type of case occurs frequently, more discussion in Appendix C. Moreover, we found that the spatial layout can be decoded back into coordinates, and inference requiring output bounding box coordinates can further alleviate the hallucination problem, more discussion in Appendix F.
@@ Appendix E Results of ChartQA
As shown in Fig. 7, the question-answer pairs in ChartQA [49] tend to involve the visual cues for reasoning. However, with only text and layout information as input, the proposed LayTextLLM inevitably have difficulties in reasoning visual-related information. Thus, on the ChartQA dataset, LayTextLLM can hardly achieve better performance than previous methods that include visual inputs. Although the visual information is not used in LayTextLLM, it can still exhibit better zero-shot ability than UniDoc [10]. After incorporating the training set of ChartQA, the performance of LayTextLLM can be boosted to 35.4%. Considering the importance of visual cues in ChartQA-like tasks, we will try to involve the visual information into LayTextLLM in future work.
@@
LayTextLLMzero (Ours) 22.8
LayTextLLMvqa (Ours) 23.4
LayTextLLMall (Ours) 35.4
@@
Appendix F Decoding Bounding Box Coordinates
We tested LayTextLLM on an in-house KIE dataset, requiring the model to output value texts along with their corresponding bounding boxes in a text format, such as "Oct 10[66,1,70,15]". We found that asking the model to output coordinates increased precision as shown in Tab. 9. This emphasizes the model’s need to search the input for corresponding information, which alleviates hallucination problems. Additionally, the model learns to combine and subtract coordinates. For example, if the output text is derived from two OCR input lines, the model will combine the corresponding OCR coordinates. Conversely, if the output is a substring of the input OCR text, the model will output the adjusted coordinates accordingly.
Model Precision Recall F-score
Output w coor 82.07 75.41 78.05
output w/o coor 83.48 74.07 77.69
Table 9: Outputting coordinates vs. not outputting coordinates